OpenStudy (anonymous):

Choose the best description of the roots of the equation 2x^2-5x+3=0. double root real and rational roots real and irrational roots imaginary roots

OpenStudy (anonymous):

help please

OpenStudy (lgbasallote):

try finding the value of \[b^2 - 4ac\]

OpenStudy (anonymous):

calculate the D which is = \[\sqrt{b ^{2}-4*a*c}\]

OpenStudy (lgbasallote):

@pulkit21384 no need for the square root

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

yeah i know, but value of discremenent is square root of b²-4ac

OpenStudy (anonymous):

@jalinnn the equation you have is of the form ax² + bx + c=0

OpenStudy (anonymous):

as you can see by comparing your equation with the general form as i have mentioned and you can get value of a=2 b=-5 c=3

OpenStudy (anonymous):

now put these values of a,b and c in the formula for d

OpenStudy (anonymous):

25=4(2*3)?

OpenStudy (anonymous):

if you mean 25-4*2*3 =1, than its ok

OpenStudy (anonymous):

now take the square root of this value which is 1

OpenStudy (anonymous):

now our conditions are D can be greater than 0 D can be smaller than 0 D can be equal to zero

OpenStudy (anonymous):

so its real and rational

OpenStudy (anonymous):

if D > 0 the roots are real and they are rational if sqare root of b² -4ac is ratioonal otherwise irrational if D<0 the roots are not real they are imaginary roots if D=0 the roots are equal

OpenStudy (anonymous):

yes you are right